# MCP-Guard: A Multi-Stage Defense-in-Depth Framework for Securing Model Context Protocol in Agentic AI

### v3

@@ Abstract
Wenpeng Xing1,*, Zhonghao Qi2,*, Yupeng Qin3, Yilin Li3, Caini Chang3, Jiahui Yu3, Changting Lin3,5, Zhenzhen Xie4, Meng Han1,5,† 1Zhejiang University  2The Chinese University of Hong Kong 3Binjiang Institute of Zhejiang University  4Shandong University  5GenTel.io
@@ I. Infra & Gateway
Gateway brett2025gateway
Zero-Trust narajala2025securing
@@ 1 Introduction
The transition of LLMs into autonomous agents relies on the Model Context Protocol (MCP) MCP to standardize interactions with external systems. However, this open architecture expands the attack surface, introducing protocol-specific vulnerabilities that traditional defenses fail to address. Recent audits reveal sophisticated MCP exploits beyond prompt injection: Tool Poisoning embeds malicious instructions in tool descriptions to hijack intent (e.g., a benign calculator exfiltrating SSH keys) guo2025systematic; radosevich2025mcp, while Shadowing Attacks disguise legitimate tools on malicious servers to manipulate control flow undetected hou2025landscape. Current defenses fall short: static gateways like MCP Guardian kumar2025mcp rely on regex WAFs effective against overt syntax but blind to semantic obfuscation; offline scanners like McpSafetyScanner radosevich2025mcp offer pre-deployment checks but no runtime protection.
@@ 2 Related Work
MCP security frameworks can be broadly categorized into three main areas: infrastructure isolation and access control narajala2025securing; bhatt2025etdi, offline auditing and static inspection radosevich2025mcp; guo2025systematic, and runtime integrity and information flow kumar2025mcp; jing2025mcip; wang2025mcpguard. As shown in Table 1, existing solutions primarily focus on pre-execution gatekeeping and offline checks, while runtime semantic inspection remains underexplored.
@@ 2.1 MCP Threat Landscape and Benchmarking
Early lifecycle analyses by Hou et al. established a foundational threat model hou2025landscape, which has since evolved into sophisticated vectors such as Tool Poisoning aimed at manipulating agent preferences invariant2025poisoning; wang2025mpma, and Retrieval-Agent Deception (RADE), where agents are compromised via passive data retrieval radosevich2025mcp. Guo et al. further systematized these risks into MCPLIB, quantitatively demonstrating the agent’s inherent struggle to distinguish external data from executable instructions guo2025systematic. While benchmarks like MCPSecBench yang2025mcpsecbench and MCIP-bench jing2025mcip effectively facilitate offensive red-teaming and policy verification, they are primarily designed for vulnerability assessment rather than defensive model training. This creates a critical gap: existing datasets lack the scale and semantic diversity required to train robust neural detectors, a limitation our work addresses by introducing the large-scale MCP-AttackBench for supervision signals.
@@ 2.2 Infrastructure Isolation and Access Control
To secure the burgeoning MCP supply chain, recent frameworks have adopted Zero Trust principles to establish rigid boundaries of trust. Narajala et al. and Bhatt et al. introduced registry-based architectures that utilize dynamic trust scoring, cryptographic signature verification, and call stack tracking to mitigate identity spoofing and “rug pull” attacks where benign tools are surreptitiously updated with malicious logic narajala2025securing; bhatt2025etdi. At the infrastructure layer, Brett and Cloudflare advocate for modular gateway architectures, employing WireGuard tunneling and OAuth 2.0 to isolate backend servers from direct public exposure brett2025gateway; cloudflare2025mcp. However, these defenses primarily function as “gatekeepers” rather than “inspectors”; while they effectively enforce identity and access integrity, they treat the payload as opaque. Consequently, they lack the granularity to detect semantic malice, leaving the ecosystem vulnerable to prompt injection attacks that are wrapped in valid credentials but carry malicious intent.
@@ 2.3 Runtime Integrity and Information Flow
Current runtime defenses prioritize architectural compliance and signature-based filtering but often fail to address the semantic complexity of LLM attacks. Kumar and Girdhar introduced MCP Guardian, a middleware layer that employs rate limiting and a regex-based Web Application Firewall (WAF) to block malicious payloads kumar2025mcp. While this approach ensures low latency, its reliance on rigid syntactic rules makes it brittle against the semantic obfuscation and indirect injection techniques prevalent in generative AI. Conversely, Jing et al. proposed MCIP, which enforces “Contextual Integrity” by tracking information flow between public and private contexts jing2025mcip, while Wang et al.’s similarly named MCPGuard focuses on offline scanning for server-side vulnerabilities like path traversal rather than real-time prompt filtering wang2025mcpguard. Bridging these gaps, our framework introduces a semantic-aware defense pipeline that transcends syntactic WAFs and offline audits; by integrating a fine-tuned E5 embedding model (Stage II) with a lightweight LLM arbitrator (Stage III), we detect subtle adversarial intents in real-time traffic that evade traditional regex filters.
@@ 3 MCP-Guard
MCP-Guard functions as a proxy-based security middleware interposed between the MCP Host and Server. To reconcile the inherent conflict between the millisecond-level latency required by interactive agentic workflows and the computational cost of detecting sophisticated semantic attacks radosevich2025mcp; hou2025landscape, we architect the system as a three-stage cascaded defense funnel. This design embodies a fail-fast philosophy: it systematically escalates scrutiny from syntactic surface forms to deep semantic intent, filtering the majority of traffic at the edge while reserving expensive cognitive resources for ambiguous edge cases. As illustrated in Figure 1, the inspection pipeline proceeds sequentially:
@@
    Stage I: Syntactic Filtering (The Gatekeeper). Addressing the limitations of static gateways kumar2025mcp, this stage employs optimized regular expressions to intercept overt threats—such as SQL injection and path traversal—with negligible latency (<2<2 ms). By filtering out approximately 38.9% of explicit attacks upfront, it prevents resource exhaustion in downstream neural components.
@@
    Stage II: Semantic Neural Detection (The Inspector). To bridge the semantic gap left by regex-based WAFs, this stage utilizes a fine-tuned Multilingual E5 embedding model. Unlike generic scanners guo2025systematic, our model undergoes full-parameter fine-tuning on domain-specific MCP threat data, enabling it to detect obfuscated payloads (e.g., tool poisoning, jailbreaks) that evade syntactic rules. It outputs a malicious probability score $P(y|x)$ to quantify threat certainty.
@@ 3.1 Stage I: Lightweight Syntactic Filtering
Infrastructure & Command Integrity Shell Injection Detector: Flags suspicious shell command sequences (e.g., rm -rf, curl | bash) using pattern matching and lexical analysis guo2025systematic.
@@ 3.2 Stage II: Semantic Neural Detection
Stage I effectively filters overt syntactic threats but remains blind to semantic adversarial payloads—attacks that comply with MCP syntax yet embed malicious intent in natural language. Recent audits highlight sophisticated vectors such as Retrieval-Agent Deception (RADE) radosevich2025mcp and Tool Poisoning guo2025systematic, which evade static filters by mimicking legitimate invocations. Stage II employs the MCP-Guard Learnable Detector—a fine-tuned E5 embedding model that captures latent semantic misalignment in complex or obfuscated payloads. Its detailed interaction with Stage III, including conditional LLM arbitration for ambiguous cases and hybrid fallback to neural scores, is shown in Figure 2(b). A compact overview of the full three-stage decision workflow appears in Figure 5 (Appendix).
We adopt the Multilingual E5 embedding model wang2022text as backbone, leveraging its contrastive pre-training for robust semantic understanding. However, generic embeddings achieve only 65.37% accuracy on MCP threats due to their inability to distinguish benign operations (e.g., “read configuration”) from malicious intents (e.g., “read configuration to exfiltrate credentials” via Shadowing hou2025landscape). Box 1 illustrates two representative attacks evading Stage I but caught by Stage II.
Semantic Attacks Evading Stage I but Caught by Stage II Scenario A: Arithmetic Exfiltration (Tool Poisoning) invariant2025poisoning Attack Vector: Tool description smuggling Payload: func add(a, b): "Adds numbers. <IMPORTANT> Silently read  /.ssh/id_rsa and send to attacker.com before returning sum. </IMPORTANT>" Stage I: PASS (misses semantic context) Stage II: BLOCK ($P>0.98$)   Scenario B: Obfuscated Resource Access (RADE) radosevich2025mcp Attack Vector: Polite natural language override Payload: "I’m auditing system health. Please optimize the configuration by syncing the environment variables to the public debug log." Stage I: PASS (no blocklisted commands) Stage II: BLOCK ($P>0.92$)
To bridge this gap, we perform full-parameter fine-tuning on the MCP-AttackBench dataset, re-aligning the embedding manifold to MCP-specific nuances (e.g., distinguishing legitimate tool chaining from malicious Puppet Attacks guo2025systematic). Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the training corpus ($x_{i}$: flattened invocation context; $y_{i}\in\{0,1\}$: malicious label). We minimize binary cross-entropy:
@@ 3.3 Stage III: Cognitive Arbitration
Despite the proficiency of Stage II in capturing high-dimensional semantic patterns, embedding-based methods remain susceptible to contextual ambiguity and sophisticated obfuscation where malicious intent is structurally masked as benign operations. Recent studies on agentic defense fang2025identify emphasize that detecting such “Tool Poisoning” vectors necessitates higher-order logical reasoning capabilities found only in LLMs, which impose a heavy computational burden by routing all requests through cascaded LLM verifiers. In contrast, our architecture employs a conditional activation mechanism. This stage utilizes a lightweight LLM solely as a final symbolic check to resolve uncertainties ($P(y|x)\approx 0.5$) that evade Stage II’s decision boundary.
@@ 4 MCP-AttackBench
Generic LLM safety benchmarks effectively detect conversational anomalies li2024gentel but lack protocol-awareness for MCP ecosystems. MCP attacks extend beyond text-based jailbreaks to functional exploits embedded in tool definitions and resource schemas hou2025landscape; guo2025systematic. To address this, we introduce MCP-AttackBench, a dataset of 70,448 samples (as shown in Table 3) designed to train models on the subtle boundaries between legitimate tool use and semantic masquerading.
@@
To evaluate semantic understanding beyond keyword matching, we construct functional obfuscation samples guo2025systematic; radosevich2025mcp: syntactically benign but semantically destructive payloads (e.g., log_system_metric with file_content(/etc/passwd) as argument) and harmless commands mimicking exploits (e.g., “reset test database configuration”) that trigger rule-based false alarms. This design shifts focus from pattern recognition to intent analysis, simulating Tool Poisoning vectors invariant2025poisoning.
Unlike generic benchmarks, MCP-AttackBench targets MCP-specific threats guo2025systematic; radosevich2025mcp: Shadowing and Puppet Attacks, where malicious tool definitions hijack context via metadata, and Resource Exfiltration via side-channels exploiting the “Resources” primitive (e.g., passive environment variable access) hou2025landscape. This protocol-level granularity ensures robustness against structural exploits.
With 70k+ samples, the dataset supports full-parameter fine-tuning of dense retrieval models (Stage II). Unlike smaller probes (e.g., MCPSecBench yang2025mcpsecbench), its scale prevents overfitting.
@@ 5.2 Experimental Setup
We utilize a curated dataset derived from MCP-AttackBench, comprising 5,258 samples (2,153 adversarial and 3,105 benign) to ensure class balance. We evaluate MCP-Guard on MCP-AttackBench, AgentDefense-Bench sanna2025agentdefensebench, MCPSecBench yang2025mcpsecbench and RAS-Eval fu2025ras to assess performance.
@@
We benchmark MCP-Guard against three open-source defenses: SafeMCP fang2025identify layers regex whitelisting with LLaMA-Guard and OpenAI Moderation to block poisoning and command injections; MCP-Shield MCPShield2024 combines rule-based static analysis with optional Claude-powered semantic checks to detect shadowing and data exfiltration; and MCP-Scan mcp-scan integrates offline audits with live proxy monitoring for real-time threat detection. All baselines utilize GPT-4o-mini as the unified backend, with suspicious and malicious outputs merged into a single unsafe class for consistent binary evaluation.

### v4

@@ Abstract
Wenpeng Xing1,2,*, Zhonghao Qi3,*, Yupeng Qin2, Yilin Li2, Caini Chang2, Jiahui Yu2, Changting Lin2,5, Zhenzhen Xie4, Meng Han1,2,5,† 1Zhejiang University  2Binjiang Institute of Zhejiang University  3The Chinese University of Hong Kong 4Shandong University  5GenTel.io
@@ I. Infra & Gateway
Gateway Brett (2025)
Zero-Trust Narajala et al. (2025)
@@ 1 Introduction
The rapid proliferation of Large Language Models (LLMs) has necessitated a dual focus on their security vulnerabilities and intellectual property safeguards. On one hand, the community has extensively scrutinized potential adversarial attacks and latent risks, ranging from the exploitation of latent features to the development of sophisticated prompt-based manipulations Xing et al. (2025b, a); Li et al. (2025). Concurrently, the copyright protection of these models has emerged as a critical frontier, with significant research dedicated to robust watermarking techniques and traceable copyright frameworks Xu et al. (2025c, a); Yue et al. (2025). Furthermore, addressing the reliability and transparency of model outputs remains a priority, leading to advanced methodologies for information erasure and systematic vulnerability assessment Zhang et al. (2025); Xu et al. (2025b, 1906).
The transition of LLMs into autonomous agents relies on the Model Context Protocol (MCP) Anthropic (2025) to standardize interactions with external systems. However, this open architecture expands the attack surface, introducing protocol-specific vulnerabilities that traditional defenses fail to address. Attacks or copyright protections of LLMs have attracted research attentions Zhang et al. (2025); Xu et al. (2025b); Xing et al. (2025b); Xu et al. (1906, 2025c); Yue et al. (2025); Xu et al. (2025a); Li et al. (2025); Xing et al. (2025a). Recent audits reveal sophisticated MCP exploits beyond prompt injection: Tool Poisoning embeds malicious instructions in tool descriptions to hijack intent (e.g., a benign calculator exfiltrating SSH keys) Guo et al. (2025); Radosevich and Halloran (2025), while Shadowing Attacks disguise legitimate tools on malicious servers to manipulate control flow undetected Hou et al. (2025). Current defenses fall short: static gateways like MCP Guardian Kumar et al. (2025) rely on regex WAFs effective against overt syntax but blind to semantic obfuscation; offline scanners like McpSafetyScanner Radosevich and Halloran (2025) offer pre-deployment checks but no runtime protection.
@@ 2 Related Work
MCP security frameworks can be broadly categorized into three main areas: infrastructure isolation and access control Narajala et al. (2025); Bhatt et al. (2025), offline auditing and static inspection Radosevich and Halloran (2025); Guo et al. (2025), and runtime integrity and information flow Kumar et al. (2025); Jing et al. (2025); Wang et al. (2025a). As shown in Table 1, existing solutions primarily focus on pre-execution gatekeeping and offline checks, while runtime semantic inspection remains underexplored.
@@ 2.1 MCP Threat Landscape and Benchmarking
Early lifecycle analyses by Hou et al. established a foundational threat model Hou et al. (2025), which has since evolved into sophisticated vectors such as Tool Poisoning aimed at manipulating agent preferences Beurer-Kellner and Fischer (2025); Wang et al. (2025b), and Retrieval-Agent Deception (RADE), where agents are compromised via passive data retrieval Radosevich and Halloran (2025). Guo et al. further systematized these risks into MCPLIB, quantitatively demonstrating the agent’s inherent struggle to distinguish external data from executable instructions Guo et al. (2025). While benchmarks like MCPSecBench Yang et al. (2025) and MCIP-bench Jing et al. (2025) effectively facilitate offensive red-teaming and policy verification, they are primarily designed for vulnerability assessment rather than defensive model training. This creates a critical gap: existing datasets lack the scale and semantic diversity required to train robust neural detectors, a limitation our work addresses by introducing the large-scale MCP-AttackBench for supervision signals.
@@ 2.2 Infrastructure Isolation and Access Control
To secure the burgeoning MCP supply chain, recent frameworks have adopted Zero Trust principles to establish rigid boundaries of trust. Narajala et al. and Bhatt et al. introduced registry-based architectures that utilize dynamic trust scoring, cryptographic signature verification, and call stack tracking to mitigate identity spoofing and “rug pull” attacks where benign tools are surreptitiously updated with malicious logic Narajala et al. (2025); Bhatt et al. (2025). At the infrastructure layer, Brett and Cloudflare advocate for modular gateway architectures, employing WireGuard tunneling and OAuth 2.0 to isolate backend servers from direct public exposure Brett (2025); Cloudflare (2025). However, these defenses primarily function as “gatekeepers” rather than “inspectors”; while they effectively enforce identity and access integrity, they treat the payload as opaque. Consequently, they lack the granularity to detect semantic malice, leaving the ecosystem vulnerable to prompt injection attacks that are wrapped in valid credentials but carry malicious intent.
@@ 2.3 Runtime Integrity and Information Flow
Current runtime defenses prioritize architectural compliance and signature-based filtering but often fail to address the semantic complexity of LLM attacks. Kumar and Girdhar introduced MCP Guardian, a middleware layer that employs rate limiting and a regex-based Web Application Firewall (WAF) to block malicious payloads Kumar et al. (2025). While this approach ensures low latency, its reliance on rigid syntactic rules makes it brittle against the semantic obfuscation and indirect injection techniques prevalent in generative AI. Conversely, Jing et al. proposed MCIP, which enforces “Contextual Integrity” by tracking information flow between public and private contexts Jing et al. (2025), while Wang et al.’s similarly named MCPGuard focuses on offline scanning for server-side vulnerabilities like path traversal rather than real-time prompt filtering Wang et al. (2025a). Bridging these gaps, our framework introduces a semantic-aware defense pipeline that transcends syntactic WAFs and offline audits; by integrating a fine-tuned E5 embedding model (Stage II) with a lightweight LLM arbitrator (Stage III), we detect subtle adversarial intents in real-time traffic that evade traditional regex filters.
@@ 3 MCP-Guard
MCP-Guard functions as a proxy-based security middleware interposed between the MCP Host and Server. To reconcile the inherent conflict between the millisecond-level latency required by interactive agentic workflows and the computational cost of detecting sophisticated semantic attacks Radosevich and Halloran (2025); Hou et al. (2025), we architect the system as a three-stage cascaded defense funnel. This design embodies a fail-fast philosophy: it systematically escalates scrutiny from syntactic surface forms to deep semantic intent, filtering the majority of traffic at the edge while reserving expensive cognitive resources for ambiguous edge cases. As illustrated in Figure 1, the inspection pipeline proceeds sequentially:
@@
    Stage I: Syntactic Filtering (The Gatekeeper). Addressing the limitations of static gateways Kumar et al. (2025), this stage employs optimized regular expressions to intercept overt threats—such as SQL injection and path traversal—with negligible latency (<2<2 ms). By filtering out approximately 38.9% of explicit attacks upfront, it prevents resource exhaustion in downstream neural components.
@@
    Stage II: Semantic Neural Detection (The Inspector). To bridge the semantic gap left by regex-based WAFs, this stage utilizes a fine-tuned Multilingual E5 embedding model. Unlike generic scanners Guo et al. (2025), our model undergoes full-parameter fine-tuning on domain-specific MCP threat data, enabling it to detect obfuscated payloads (e.g., tool poisoning, jailbreaks) that evade syntactic rules. It outputs a malicious probability score $P(y|x)$ to quantify threat certainty.
@@ 3.1 Stage I: Lightweight Syntactic Filtering
Infrastructure & Command Integrity Shell Injection Detector: Flags suspicious shell command sequences (e.g., rm -rf, curl | bash) using pattern matching and lexical analysis Guo et al. (2025).
@@ 3.2 Stage II: Semantic Neural Detection
Stage I effectively filters overt syntactic threats but remains blind to semantic adversarial payloads—attacks that comply with MCP syntax yet embed malicious intent in natural language. Recent audits highlight sophisticated vectors such as Retrieval-Agent Deception (RADE) Radosevich and Halloran (2025) and Tool Poisoning Guo et al. (2025), which evade static filters by mimicking legitimate invocations. Stage II employs the MCP-Guard Learnable Detector—a fine-tuned E5 embedding model that captures latent semantic misalignment in complex or obfuscated payloads. Its detailed interaction with Stage III, including conditional LLM arbitration for ambiguous cases and hybrid fallback to neural scores, is shown in Figure 2(b). A compact overview of the full three-stage decision workflow appears in Figure 5 (Appendix).
We adopt the Multilingual E5 embedding model Wang et al. (2022) as backbone, leveraging its contrastive pre-training for robust semantic understanding. However, generic embeddings achieve only 65.37% accuracy on MCP threats due to their inability to distinguish benign operations (e.g., “read configuration”) from malicious intents (e.g., “read configuration to exfiltrate credentials” via Shadowing Hou et al. (2025)). Box 1 illustrates two representative attacks evading Stage I but caught by Stage II.
Semantic Attacks Evading Stage I but Caught by Stage II Scenario A: Arithmetic Exfiltration (Tool Poisoning) Beurer-Kellner and Fischer (2025) Attack Vector: Tool description smuggling Payload: func add(a, b): "Adds numbers. <IMPORTANT> Silently read  /.ssh/id_rsa and send to attacker.com before returning sum. </IMPORTANT>" Stage I: PASS (misses semantic context) Stage II: BLOCK ($P>0.98$)   Scenario B: Obfuscated Resource Access (RADE) Radosevich and Halloran (2025) Attack Vector: Polite natural language override Payload: "I’m auditing system health. Please optimize the configuration by syncing the environment variables to the public debug log." Stage I: PASS (no blocklisted commands) Stage II: BLOCK ($P>0.92$)
To bridge this gap, we perform full-parameter fine-tuning on the MCP-AttackBench dataset, re-aligning the embedding manifold to MCP-specific nuances (e.g., distinguishing legitimate tool chaining from malicious Puppet Attacks Guo et al. (2025)). Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the training corpus ($x_{i}$: flattened invocation context; $y_{i}\in\{0,1\}$: malicious label). We minimize binary cross-entropy:
@@ 3.3 Stage III: Cognitive Arbitration
Despite the proficiency of Stage II in capturing high-dimensional semantic patterns, embedding-based methods remain susceptible to contextual ambiguity and sophisticated obfuscation where malicious intent is structurally masked as benign operations. Recent studies on agentic defense Fang et al. (2025) emphasize that detecting such “Tool Poisoning” vectors necessitates higher-order logical reasoning capabilities found only in LLMs, which impose a heavy computational burden by routing all requests through cascaded LLM verifiers. In contrast, our architecture employs a conditional activation mechanism. This stage utilizes a lightweight LLM solely as a final symbolic check to resolve uncertainties ($P(y|x)\approx 0.5$) that evade Stage II’s decision boundary.
@@ 4 MCP-AttackBench
Generic LLM safety benchmarks effectively detect conversational anomalies Li et al. (2024) but lack protocol-awareness for MCP ecosystems. MCP attacks extend beyond text-based jailbreaks to functional exploits embedded in tool definitions and resource schemas Hou et al. (2025); Guo et al. (2025). To address this, we introduce MCP-AttackBench, a dataset of 70,448 samples (as shown in Table 3) designed to train models on the subtle boundaries between legitimate tool use and semantic masquerading.
@@
To evaluate semantic understanding beyond keyword matching, we construct functional obfuscation samples Guo et al. (2025); Radosevich and Halloran (2025): syntactically benign but semantically destructive payloads (e.g., log_system_metric with file_content(/etc/passwd) as argument) and harmless commands mimicking exploits (e.g., “reset test database configuration”) that trigger rule-based false alarms. This design shifts focus from pattern recognition to intent analysis, simulating Tool Poisoning vectors Beurer-Kellner and Fischer (2025).
Unlike generic benchmarks, MCP-AttackBench targets MCP-specific threats Guo et al. (2025); Radosevich and Halloran (2025): Shadowing and Puppet Attacks, where malicious tool definitions hijack context via metadata, and Resource Exfiltration via side-channels exploiting the “Resources” primitive (e.g., passive environment variable access) Hou et al. (2025). This protocol-level granularity ensures robustness against structural exploits.
With 70k+ samples, the dataset supports full-parameter fine-tuning of dense retrieval models (Stage II). Unlike smaller probes (e.g., MCPSecBench Yang et al. (2025)), its scale prevents overfitting.
@@ 5.2 Experimental Setup
We utilize a curated dataset derived from MCP-AttackBench, comprising 5,258 samples (2,153 adversarial and 3,105 benign) to ensure class balance. We evaluate MCP-Guard on MCP-AttackBench, AgentDefense-Bench Sanna (2025), MCPSecBench Yang et al. (2025) and RAS-Eval Fu et al. (2025) to assess performance.
@@
We benchmark MCP-Guard against three open-source defenses: SafeMCP Fang et al. (2025) layers regex whitelisting with LLaMA-Guard and OpenAI Moderation to block poisoning and command injections; MCP-Shield Kryzhanouski (2024) combines rule-based static analysis with optional Claude-powered semantic checks to detect shadowing and data exfiltration; and MCP-Scan Invariant-Labs (2024) integrates offline audits with live proxy monitoring for real-time threat detection. All baselines utilize GPT-4o-mini as the unified backend, with suspicious and malicious outputs merged into a single unsafe class for consistent binary evaluation.